# Uncertainty Quantification for LLM-Based Survey Simulations

### v1

@@ 3.1 Problem Formulation
We consider constructing the confidence set by using simulated responses from an LLM. Given a profile $z$, a survey question $\psi$ and a prompt $p$, the LLM simulates a response $y^{\mathsf{syn}}$ from a distribution $\mathcal{Q}^{\mathsf{syn}}(~{}\cdot\mid z,\psi,p)$ which aims to mimic $\mathcal{Q}(~{}\cdot\mid z,\psi)$. We can generate synthetic profiles $\{z^{\mathsf{syn}}_{i}\}_{i=1}^{K}$ from some distribution $\mathcal{P}^{\mathsf{syn}}$, then feed them into the LLM along with $\psi$ and $p$. The LLM then generates synthetic responses $\{y^{\mathsf{syn}}_{i}\}_{i=1}^{K}$, where $y^{\mathsf{syn}}_{i}\sim\mathcal{Q}^{\mathsf{syn}}(~{}\cdot\mid z^{\mathsf{syn}}_{i},\psi,p)$. Here $K$ is the simulation budget.
@@
We also simulate LLM responses to these $m$ survey questions. For each $j\in[m]$, we feed synthetic profiles $\{z^{\mathsf{syn}}_{j,i}\}_{i=1}^{K}$, the question $\psi_{j}$ and the prompt $p$ into the LLM, which simulates responses $\mathcal{D}^{\mathsf{syn}}_{j}=\{y^{\mathsf{syn}}_{j,i}\}_{i=1}^{K}$ with $y^{\mathsf{syn}}_{j,i}\sim\mathcal{Q}^{\mathsf{syn}}(~{}\cdot\mid z^{\mathsf{syn}}_{j,i},\psi_{j},p)$. The datasets $\{\mathcal{D}_{j}\}_{j=1}^{m}$ and $\{\mathcal{D}^{\mathsf{syn}}_{j}\}_{j=1}^{m}$ will be used to select a confidence set from $\{{\mathcal{S}}^{\mathsf{syn}}(k)\}_{k=1}^{K}$. We make the same Assumptions 2.1 and 2.2 as in Section 2.
@@
Problem 1 (Uncertainty quantification).
@@ 3.2 General Methodology for Sample Size Selection
We now present our general methodology for Problem 1. For each $j\in[m]$, we form confidence sets similar to (3.1) using the synthetic data $\mathcal{D}^{\mathsf{syn}}_{j}$:
@@
Ideally, we would like to pick $k$ such that $(1-\alpha)$ coverage is achieved empirically over the $m$ survey functions:
@@
as a proxy for the empirical miscoverage. Since $\theta(\psi_{j})\in{\mathcal{S}}_{j}$ happens with probability $\gamma$, then the frequency of having ${\mathcal{S}}_{j}\not\subseteq{\mathcal{S}}^{\mathsf{syn}}_{j}(k)$ is at least $\gamma$ times the frequency of $\theta(\psi_{j})\not\in{\mathcal{S}}^{\mathsf{syn}}_{j}(k)$. Roughly speaking,
@@
It is worth noting that our method achieves this coverage without any assumptions on the qualities of the LLM and the procedure $\mathcal{C}$ for confidence set construction. Nevertheless, the size of the chosen confidence set ${\mathcal{S}}^{\mathsf{syn}}(\widehat{k})$, in terms of the true coverage rate and size, depend on these factors. If there is a large alignment gap between the LLM and the human population, then ${\mathcal{S}}^{\mathsf{syn}}(\widehat{k})$ will inevitably be large.
@@ 4.1 Experiment Setup
We consider $8$ LLMs: GPT-3.5-Turbo (gpt-3.5-turbo), GTP-4o (gpt-4o), and GPT-4o-mini (gpt-4o-mini) (OpenAI, 2022, 2024b, 2024a); Claude 3.5 Haiku (claude-3-5-haiku-20241022) Anthropic (2024); Llama 3.1 8B (Llama-3-8B-Instruct-Turbo) and Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo) (Dubey et al., 2024); Mistral 7B (Mistral-7B-Instruct-v0.3) (Jiang et al., 2023); DeepSeek-V3 (DeepSeek-V3) (Liu et al., 2024).
@@
We use two datasets for survey questions, each corresponding to one uncertainty quantification task. The first dataset is the OpinionQA dataset created by Santurkar et al. (2023). It was built from Pew Research’s American Trends Panel [^1] https://www.pewresearch.org/the-american-trends-panel/, and contains the general US population’s responses to survey questions spanning topics such as science, politics, and health. After pre-processing we have 385 unique questions and 1,476,868 responses to these questions from at least 32,864 people. These questions have $5$ choices corresponding to ordered sentiments which we map to sentiment scores $-1,-\frac{1}{3},0,\frac{1}{3},1$. Each question has at least $400$ responses. For each response, we have information on their political profile, religious affiliation, educational background, socio-economic status, etc. This information is used as their profiles to generate synthetic profiles. See Section C.1 for more information on this dataset, including example questions, profile features, and the generation of synthetic profiles and answers. We consider the task of constructing a confidence interval for the US population’s average sentiment score for a survey question. This is the setup in Example A.3.
The second dataset is the EEDI dataset created by He-Yueya et al. (2024), which was built upon the NeurIPS 2020 Education Challenge dataset (Wang et al., 2021). It consists of students’ responses to mathematics multiple-choice questions on the Eedi online educational platform [^2] https://eedi.com/. The dataset contains 573 unique questions and 443,433 responses to these questions from 2,287 students. All questions have four choices (A, B, C, D). Out of these questions, we use questions that have at least $100$ student responses. Excluding questions with graphs or diagrams, we are left with a total of $412$ questions. For each student, we have information on their gender, age, and socioeconomic status. This information is used as their profiles to generate synthetic profiles. See Section C.2 for more information on this dataset, including example questions, profile distribution, and the generation of synthetic profiles and answers. We consider the task of constructing a confidence interval for the probability of a student answering a question correctly. This is similar to the setup in Section 2 and Example 3.1.
@@ 4.2 Experiment Procedure
where $\bar{y}_{j}=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}y_{j,i}$ and $s_{j}=\sqrt{\bar{y}_{j}(1-\bar{y}_{j})}$. Since $n_{j}$ is at least $100$, ${\mathcal{S}}_{j}$ has approximately $\gamma$ coverage probability. In Figure 5 of Appendix D, we provide a visualization for the selection of $\widehat{k}$.
@@
We use $\mathscr{D}^{\mathrm{te}}$ to evaluate the quality of the chosen simulation sample size $\widehat{k}$. As the true population mean $\theta(\psi)$ is unavailable, the true coverage probability $\mathbb{P}\big{(}\theta(\psi)\in{\mathcal{S}}^{\mathsf{syn}}(\widehat{k})\big{)}$ cannot be computed. However, we can apply the same idea as (3.5) in Section 3 to compute a proxy for the miscoverage level. For each survey question $j\in\mathcal{J}_{2}$, the selected sample size $\widehat{k}$ leads to the synthetic confidence set ${\mathcal{S}}^{\mathsf{syn}}_{j}(\widehat{k})=\mathcal{C}\big{(}\{y^{\mathsf{syn}}_{j,i}\}_{i=1}^{\widehat{k}}\big{)}$. We form the confidence set ${\mathcal{S}}_{j}$ from real data $\mathcal{D}_{j}$ as in (4.2) and define
$\widetilde{L}(\widehat{k})=\frac{1}{\gamma}\cdot\frac{1}{|\mathcal{J}_{\mathrm{te}}|}\sum_{j\in\mathcal{J}_{\mathrm{te}}}\operatorname{\mathbf{1}}\{{\mathcal{S}}_{j}\not\subseteq{\mathcal{S}}^{\mathsf{syn}}_{j}(\widehat{k})\}.$ (4.3)
@@
$\gamma\cdot\mathbb{P}\big{(}\theta(\psi)\not\in{\mathcal{S}}^{\mathsf{syn}}(k)\big{)}\leq\mathbb{P}\big{(}{\mathcal{S}}_{j}\not\subseteq{\mathcal{S}}^{\mathsf{syn}}_{j}(k)\big{)}=\gamma\cdot\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}.$
Thus, if $\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}\leq\alpha$, then $\mathbb{P}\big{(}\theta(\psi)\not\in{\mathcal{S}}^{\mathsf{syn}}(\widehat{k})\big{)}\leq\alpha$ must hold. To that end, we will test a hypothesis $H_{0}:\;\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}\leq\alpha$ against its alternative $H_{1}:\;\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}>\alpha$.
@@ 4.3 Experiment Results
Figure 3: Histograms of $p$-values for the hypothesis test $\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}\leq\alpha$ against $\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}>\alpha$ across various LLMs and $\alpha$’s over the OpinionQA (left) and EEDI (right) datasets.
In Figure 4, we plot the average $\widehat{k}$ over the $100$ random splits for various LLMs on the OpinionQA and EEDI datasets. The error bars represent 95% confidence intervals. In general, a larger $\widehat{k}$ means that the LLM has stronger simulation power. On the OpinionQA dataset, GPT-4o has the best performance. On the EEDI dataset, DeepSeek-V3 has the best performance. Interestingly, on the OpinionQA dataset all LLMs clearly outperform the random benchmark, while on the EEDI dataset only DeepSeek-V3 and GPT-4o seem to outperform the random benchmark. Moreover, LLMs exhibit uniformly higher $\widehat{k}$ on the OpinionQA dataset than on the EEDI dataset, suggesting higher fidelity in simulating subjective opinions than in simulating answers to mathematics questions.
Figure 4: Average $\widehat{k}$ for various LLMs and $\alpha$ over the OpinionQA (left) and EEDI (right) datasets.
The experiment results demonstrate the importance of a disciplined approach to using synthetic samples. The ease of LLM-based simulation makes it tempting to generate a large number of responses per question. However, as can be seen from the figures, there is great heterogeneity in the simulation power of different LLMs over different datasets: the largest $\widehat{k}$ is below 100, while the smallest $\widehat{k}$ could be in the single digits. This means that there is real peril in using too many synthetic samples and being overly confident in the results.
@@ 5 Discussions
We developed a general approach for converting imperfect LLM-based survey simulations into statistically valid confidence sets for population statistics of human responses. It identifies a simulation sample size which is useful for future simulation tasks and which reveals the degree of misalignment between the LLM and the target human population.
Several future directions are worth exploring. First, our approach does not explicitly minimize the size of the prediction set. A natural question is whether we can incorporate a size minimization procedure to produce smaller confidence sets with good coverage. Second, it would be interesting to see if our approach can be combined with debiasing methods to give more informative confidence sets. Finally, as prompt engineering is known to have crucial effects on the quality of LLM generations, it is worth investigating the impacts of prompts on the selected simulation sample size $\widehat{k}$, and how prompt engineering can be leveraged to reduce the misalignment gap between LLM simulations and true human responses.
@@ Acknowledgement
Appendix A More Examples for Section 3
In this section, we provide more examples for the general problem framework in Section 3. In these examples, the survey responses can be real-valued or multi-dimensional.
@@ Appendix A More Examples for Section 3
Suppose an organization is interested in performing a public survey in a city. Each survey question $\psi$ is a multiple-choice question with $5$ options. An example is “How often do you talk to your neighbors?”, with $5$ choices “Basically every day”, “A few times a week”, “A few times a month”, “Once a month”, and “Less than once a month”. Every $z\in\mathcal{Z}$ is a person’s profile (e.g., age, gender, occupation), the response space $\mathcal{Y}$ is the standard orthonormal basis $\{e_{i}\}_{i=1}^{5}$ in $\mathbb{R}^{5}$, where $y=e_{i}$ indicates that a person chooses the $i$-th option. We can take $\theta(\psi)=\mathbb{E}_{y\sim\mathcal{Q}(\cdot\mid\psi)}[y]\in\mathbb{R}^{5}$, which summarizes the proportion of people that choose the $i$-th option. An LLM can be used to simulate people’s answers to the survey question.
@@
Consider the setup in Example A.2. When the $5$ choices in a survey question correspond to ordered sentiments, we can map them to numeric scores, say, $v=(-1,-\frac{1}{3},0,\frac{1}{3},1)^{\top}$. Then the statistic $\widetilde{\theta}(\psi)=\langle v,\theta(\psi)\rangle$ reflects the population’s average sentiment in the survey question $\psi$.

### v3

@@ 3.1 Problem Formulation
We consider constructing the confidence set by using simulated responses from an LLM. Given a profile $z$, a survey question $\psi$ and a prompt $p$, the LLM simulates a response $y^{\mathsf{syn}}$ from a distribution $\mathcal{Q}^{\mathsf{syn}}(~{}\cdot\mid z,\psi,p)$ which aims to mimic $\mathcal{Q}(~{}\cdot\mid z,\psi)$. We can generate i.i.d. synthetic profiles $\{z^{\mathsf{syn}}_{i}\}_{i=1}^{K}\sim\mathcal{P}^{\mathsf{syn}}$, then feed them into the LLM along with $\psi$ and $p$. The LLM then generates synthetic responses $\{y^{\mathsf{syn}}_{i}\}_{i=1}^{K}$, where $y^{\mathsf{syn}}_{i}\sim\mathcal{Q}^{\mathsf{syn}}(~{}\cdot\mid z^{\mathsf{syn}}_{i},\psi,p)$. Here $K\in\mathbb{Z}_{+}$ is the simulation budget.
@@
We also simulate LLM responses to these $m$ survey questions. For each $j\in[m]$, we feed i.i.d. synthetic profiles $\{z^{\mathsf{syn}}_{j,i}\}_{i=1}^{K}\sim\mathcal{P}^{\mathsf{syn}}$, the question $\psi_{j}$ and the prompt $p$ into the LLM, which then simulates responses $\mathcal{D}^{\mathsf{syn}}_{j}=\{y^{\mathsf{syn}}_{j,i}\}_{i=1}^{K}$ with $y^{\mathsf{syn}}_{j,i}\sim\mathcal{Q}^{\mathsf{syn}}(~{}\cdot\mid z^{\mathsf{syn}}_{j,i},\psi_{j},p)$. The datasets $\{\mathcal{D}_{j}\}_{j=1}^{m}$ and $\{\mathcal{D}^{\mathsf{syn}}_{j}\}_{j=1}^{m}$ will be used to select a confidence set from $\{{\mathcal{S}}^{\mathsf{syn}}(k)\}_{k=1}^{K}$. We make the same Assumptions 2.1 and 2.2 as in Section 2.
@@
Problem 3.1 (Uncertainty quantification).
@@ 3.2 General Methodology for Sample Size Selection
We now present our general methodology for Problem 3.1. For each $j\in[m]$, we form confidence sets similar to (3.1) using the synthetic data $\mathcal{D}^{\mathsf{syn}}_{j}$:
@@
Ideally, we would like to pick $k$ such that $(1-\alpha)$ coverage is achieved empirically over the $m$ survey questions:
@@
as a proxy for the empirical miscoverage. Since $\theta(\psi_{j})\in{\mathcal{S}}_{j}$ with probability $\gamma$, then the frequency of having ${\mathcal{S}}_{j}\not\subseteq{\mathcal{S}}^{\mathsf{syn}}_{j}(k)$ is at least $\gamma$ times the frequency of $\theta(\psi_{j})\not\in{\mathcal{S}}^{\mathsf{syn}}_{j}(k)$. Roughly speaking,
@@
It is worth noting that our method achieves this coverage without any assumptions on the qualities of the LLM and the procedure $\mathcal{C}$ for confidence set construction. Nevertheless, the size of the chosen confidence set ${\mathcal{S}}^{\mathsf{syn}}(\widehat{k})$ depends on these factors. If there is a large alignment gap between the LLM and the human population, then ${\mathcal{S}}^{\mathsf{syn}}(\widehat{k})$ will inevitably be large.
@@ 4.1 Experiment Setup
We consider $8$ LLMs: GPT-3.5-Turbo (gpt-3.5-turbo), GPT-4o (gpt-4o), and GPT-4o-mini (gpt-4o-mini) (OpenAI, 2022, 2024b, 2024a); Claude 3.5 Haiku (claude-3-5-haiku-20241022) Anthropic (2024); Llama 3.1 8B (Llama-3-8B-Instruct-Turbo) and Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo) (Dubey et al., 2024); Mistral 7B (Mistral-7B-Instruct-v0.3) (Jiang et al., 2023); DeepSeek-V3 (DeepSeek-V3) (Liu et al., 2024).
@@
We use two datasets for survey questions, each corresponding to one uncertainty quantification task. The first dataset is the OpinionQA dataset Santurkar et al. (2023). It was built from Pew Research’s American Trends Panel [^1] https://www.pewresearch.org/the-american-trends-panel/, and contains the general US population’s responses to survey questions spanning topics such as science, politics, and health. After pre-processing we have 385 unique questions and 1,476,868 responses to these questions from at least 32,864 people. These questions have $5$ choices corresponding to ordered sentiments which we map to sentiment scores $-1,-\frac{1}{3},0,\frac{1}{3},1$. Each question has at least $400$ responses. For each response, we have information on the surveyee’s political profile, religious affiliation, educational background, socio-economic status, etc., which we use to generate synthetic profiles. See Section C.1 for more details. We consider the task of constructing a confidence interval for the US population’s average sentiment score for a survey question. This is the setup in Example A.3.
The second dataset is the EEDI dataset created by He-Yueya et al. (2024), which was built upon the NeurIPS 2020 Education Challenge dataset (Wang et al., 2021). It consists of students’ responses to mathematics multiple-choice questions on the Eedi online educational platform [^2] https://eedi.com/. The dataset contains 573 unique questions and 443,433 responses to these questions from 2,287 students. All questions have four choices (A, B, C, D). Out of these questions, we use questions that have at least $100$ student responses. Excluding questions with graphs or diagrams, we are left with a total of $412$ questions. For each student, we have information on their gender, age, and socioeconomic status, which we use to generate synthetic profiles. See Section C.2 for more details. We consider the task of constructing a confidence interval for the probability of a student answering a question correctly. This is similar to the setup in Example 3.1.
@@ 4.2 Experiment Procedure
where $\bar{y}_{j}=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}y_{j,i}$ and $s_{j}=\sqrt{\bar{y}_{j}(1-\bar{y}_{j})}$. Since $n_{j}$ is at least $100$, ${\mathcal{S}}_{j}$ has approximately $\gamma$ coverage probability. In Figure 5 of Section C.3, we provide a visualization for the selection of $\widehat{k}$.
@@
We use $\mathscr{D}^{\mathrm{te}}$ to evaluate the quality of the chosen simulation sample size $\widehat{k}$. As the true population parameter $\theta(\psi)$ is unavailable, the true coverage probability $\mathbb{P}\big{(}\theta(\psi)\in{\mathcal{S}}^{\mathsf{syn}}(\widehat{k})\big{)}$ cannot be computed. However, we can apply the same idea as (3.5) in Section 3 to compute a proxy for the miscoverage level. For each survey question $j\in\mathcal{J}_{\mathrm{te}}$, the selected sample size $\widehat{k}$ leads to the synthetic confidence set ${\mathcal{S}}^{\mathsf{syn}}_{j}(\widehat{k})=\mathcal{C}\big{(}\{y^{\mathsf{syn}}_{j,i}\}_{i=1}^{\widehat{k}}\big{)}$. We form the confidence set ${\mathcal{S}}_{j}$ from real data $\mathcal{D}_{j}$ as in (4.2) and define
$\widetilde{L}(k)=\frac{1}{\gamma}\cdot\frac{1}{|\mathcal{J}_{\mathrm{te}}|}\sum_{j\in\mathcal{J}_{\mathrm{te}}}\operatorname{\mathbf{1}}\big{\{}{\mathcal{S}}_{j}\not\subseteq{\mathcal{S}}^{\mathsf{syn}}_{j}(k)\big{\}}.$ (4.3)
@@
$\gamma\cdot\mathbb{P}\big{(}\theta(\psi)\not\in{\mathcal{S}}^{\mathsf{syn}}(k)\big{)}\leq\mathbb{P}\big{(}{\mathcal{S}}_{j}\not\subseteq{\mathcal{S}}^{\mathsf{syn}}_{j}(k)\big{)}=\gamma\cdot\mathbb{E}\big{[}\widetilde{L}(k)\big{]}.$
Thus, if $\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}\leq\alpha$, then $\mathbb{P}\big{(}\theta(\psi)\not\in{\mathcal{S}}^{\mathsf{syn}}(\widehat{k})\big{)}\leq\alpha$ must hold. To this end, we will test a hypothesis $H_{0}:\;\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}\leq\alpha$ against its alternative $H_{1}:\;\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}>\alpha$.
@@ 4.3 Experiment Results
Figure 2: Histograms of $p$-values for the hypothesis test $\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}\leq\alpha$ against $\mathbb{E}\big{[}\widetilde{L}(\widehat{k})\big{]}>\alpha$ across various LLMs and $\alpha$’s over the OpinionQA dataset (left) and the EEDI dataset (right).
Selected simulation sample size.
In Figure 3, we plot the average $\widehat{k}$ over the $100$ random splits for various LLMs on the OpinionQA and EEDI datasets, respectively. The error bars represent 95% confidence intervals.
Figure 3: Average $\widehat{k}$ for various LLMs and $\alpha$ over the OpinionQA dataset (left) and the EEDI dataset (right).
In general, a larger $\widehat{k}$ means that the LLM has a stronger simulation power. On the OpinionQA dataset, GPT-4o has the best performance. On the EEDI dataset, DeepSeek-V3 has the best performance, followed by GPT-4o and Claude 3.5 Haiku. Interestingly, on the OpinionQA dataset all LLMs clearly outperform the random benchmark, while on the EEDI dataset only DeepSeek-V3 and GPT-4o seem to outperform the random benchmark. Moreover, LLMs exhibit uniformly higher $\widehat{k}$ on the OpinionQA dataset than on the EEDI dataset, suggesting higher fidelity in simulating subjective opinions to social problems than in simulating student answers to mathematics questions.
The experiment results demonstrate the importance of a disciplined approach to using synthetic samples. The ease of LLM-based simulation makes it tempting to generate a large number of responses per question. However, our results show that there is great heterogeneity in the simulation powers of different LLMs over different datasets: the largest $\widehat{k}$ is below 100, while the smallest $\widehat{k}$ could be in the single digits. This means that there is real peril in using too many synthetic samples and being overly confident in the results.
@@ 5 Discussions
We developed a general approach for converting imperfect LLM-based survey simulations into statistically valid confidence sets for population parameters of human responses. It identifies a simulation sample size which is useful for future simulation tasks and which quantifies the fidelity of LLM simulations. Numerical experiments on real datasets verified the coverage guarantees of our approach, and revealed that existing LLMs exhibited higher fidelity in simulating opinions to social problems than in simulating student answers to mathematics questions.
Several future directions are worth exploring. First, our approach does not explicitly minimize the size of the prediction set. A natural question is whether we can incorporate a size minimization procedure to produce smaller confidence sets with good coverage. Second, it would be interesting to see if our approach can be combined with debiasing methods to give more informative confidence sets. Finally, as prompt engineering is known to have crucial effects on the quality of LLM generations, it is worth investigating the impacts of prompts on the selected simulation sample size $\widehat{k}$, and how prompt engineering can be leveraged to improve the fidelity of LLM simulations.
@@ Acknowledgement
Appendix A Illustrations and Examples
We first present a visualization of the relation between the simulation sample size and the LLM’s simulation power in Figure 4.
Figure 4: An interpretation of the simulation sample size $\widehat{k}$ as the size of human population that an LLM can represent. Generating outputs from the LLM can be thought of as “resampling” from $\widehat{k}$ human samples that make up the LLM. The figure is generated by DALL-E (Ramesh et al., 2021), and borrows ideas from the Mechanical Turk, a chess-playing machine from the 18th century with a human player hidden inside.
Below we provide more examples for the general problem framework in Section 3. The survey responses can be real-valued or multi-dimensional.
@@ Appendix A Illustrations and Examples
Suppose an organization is interested in performing a public survey in a city. Each survey question $\psi$ is a multiple-choice question with $5$ options. An example is “How often do you talk to your neighbors?”, with $5$ choices “Basically every day”, “A few times a week”, “A few times a month”, “Once a month”, and “Less than once a month”. Every $z\in\mathcal{Z}$ is a person’s profile (e.g., age, gender, occupation), the response space $\mathcal{Y}$ is the standard orthonormal basis $\{e_{i}\}_{i=1}^{5}$ in $\mathbb{R}^{5}$, where $y=e_{i}$ indicates that a person chooses the $i$-th option. We can take $\theta(\psi)=\mathbb{E}_{y\sim\mathcal{Q}(\cdot\mid\psi)}[y]\in\mathbb{R}^{5}$, which summarizes the proportion of people that choose each option. An LLM can be used to simulate people’s answers to the survey question.
@@
Consider the setup in Example A.2. When the $5$ choices in a survey question correspond to ordered sentiments, we can map them to numeric scores, say, $v=(-1,-\frac{1}{3},0,\frac{1}{3},1)^{\top}$. Then the statistic $\widetilde{\theta}(\psi)=\langle v,\theta(\psi)\rangle$ reflects the population’s average sentiment for the survey question $\psi$.